OpenStudy (anonymous):

how do you do integral on x^3-2x-x

OpenStudy (anonymous):

When you have a sum or a difference under the sign of an integral, you can split it into many integrals. By doing so, you will have three different integrals, but they will be easy to calculate.

OpenStudy (anonymous):

1/4x^4- 2/2x^2-1/2x^2 is that right?

OpenStudy (anonymous):

Yes, this is correct, and of course you need to add the arbitrary constant C to your final answer.

OpenStudy (anonymous):

+ C o ok and can it be simplyfied in any way?

OpenStudy (anonymous):

well, 2/2=1 and you could factor out x^2.

OpenStudy (anonymous):

You should simply x^3 - 3x prior to taking integral!

OpenStudy (anonymous):

i have 1/4 X^4-3/2 x^2 in my textbook as an answer I don't see how they got there

OpenStudy (anonymous):

i have no clue what you did there

OpenStudy (anonymous):

no how did you get there

OpenStudy (anonymous):

1/4x^4- 2/2x^2-1/2x^2 < how does that get to x³- 3x

OpenStudy (anonymous):

1/4 X^4-3/2 x^2 is the answer i have in my book

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

ah!

OpenStudy (anonymous):

∫ x³ dx = x^4/4 - 3∫ xdx = -3x²/2

OpenStudy (anonymous):

-2x-x= -3x

OpenStudy (anonymous):

i see it now very simple

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

Yup, indeed it issss!